# On Topology and Geometry of Hessian Manifolds

### v2

@@ Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is termed an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is called a Hessian metric on affine manifold $(M,\nabla)$, if $g$ can be expressed locally as the Hessian quadratic form of some smooth potential $f$. An affine manifold equipped with a Hessian metric is termed a Hessian manifold. We study the topology and geometry of Hessian manifolds.
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and the proof that its fundamental group must be infinite and torsion-free. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the canonical and obstruction bundles, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian immersions, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature, and classify all complete Hessian surfaces.
@@
Hessian manifold, Affine differential geometry, Flat line bundle, Legendre transformation, Lagrangian immersion.
@@ 1.1 Introduction and Background
A central achievement of our work is to reveal that Hessian metrics of Koszul-type are part of a much larger network of connections than previously understood. Whilst the work of Shima [15] established an important equivalence between the existence of a Koszul type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Corollary2.48) and admitting a space-like Lagrangian immersion (Corollary3.24) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
@@
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem2.43), revealing a remarkably rigid underlying structure. This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition3.41).
@@
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem4.5). We also derive severe constraints on the intersection form for Hessian 4-manifolds, and then prove that any closed orientable Hessian-Einstein manifold of dimension four must be flat (Proposition4.20). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
@@
This paper is organized as follows:
Section1 reviews fundamental definitions and introduces a warped product construction as first examples. Section2 suggests key topological obstructions, concerning the vanishing of Euler characteristic and the infinitude of fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section3 establishes the generalised Legendre transform, and a central equivalence between Hessian metrics of Koszul type and several other geometric structures, including positive flat line bundles and space-like Lagrangian immersions. Curvature properties and link to complex analysis are also discussed. Finally, Section4 applies these results to low dimensions, yielding a complete topological classification of Hessian surfaces and deriving strong constraints for Hessian 4-manifolds. An appendix is included to support our classification of compact Hessian manifolds with Abelian fundamental groups.
@@ 1.3 Warped Product Construction
Now we state the fundamental construction theorem of warped products over a Hessian metric.
@@ 2.2 Fundamental Group and Euler Number
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n$. If the fundamental group of $M$ is Abelian, then $M$ is homeomorphic to the flat $n$-torus $\mathbb{T}^{n}$.
@@ 2.3.1 The Picard group
Since the connection $D$ is flat, its connection 1-forms are locally exact. Without lose of generality, we may assume that there exists $f\in C^{\infty}(U_{\alpha})$ such that the restriction of $D$ on $U_{\alpha}$ is $d+df$.
@@
Now we introduce the isomorphism classes in the category of flat $\mathbb{R}$-line bundle over $M$.
@@ 2.3.2 The canonical line bundle
For an oerientable Riemannian manifold $(M,h)$, we denote by $\operatorname{vol}_{h}$ its volume form.
@@
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then $M$ is a mapping torus.
@@
Let $\eta\in\Omega^{1}(M)$ be a $1$-form potential of $((M,\nabla),h)$. View $\eta\colon M\rightarrow T^{*}M$ as an immersion. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
@@
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or $M$ is a mapping torus.
@@
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem2.38. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition2.37, we have that $\varkappa$ is nowhere vanishing by parallel transport. Therefore by Theorem2.41 we conclude that $M$ is a mapping torus. ∎
@@ 3.1 Digression on Complex Geometry
Now, we are ready to state and prove the main theorem of this section.
@@
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper real analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
@@ 3.2 Legendre Transformation
The following proposition characterizes the universal coverings of compact radiant affine manifolds as open cones in the Euclidean spaces.
@@
See [11] for a proof of the fact that the conjugate connection is well-defined for every connection on the tangent bundle of a pseudo-Riemannian manifold.
@@
3.3 Space-like Lagrangian Immersions
In this section, we characterize Hessian manifolds of Koszul type as space-like Lagrangians.
@@ 3.3 Space-like Lagrangian Immersions
Note that, the Riemann extension of a flat torsion-free connection is of neutral signature.
@@
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$, a $1$-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian immersion of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an immersed Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the Riemann extension of $\nabla$.
The following theorem shows that, in the category of affine manifolds, space-like immersed Lagrangians are precisely those Hessian manifolds of Koszul type.
@@
    $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$;
@@
Recall from differential topology that a smooth cross-section of a fiber bundle is always an immersion.
Denote by $\tau$ the tautological 1-form of $T^{*}M$, and denote by $\omega$ the standard symplectic form on $T^{*}M$. Then we have $\omega=-d\tau$ and $\eta^{*}\tau=\eta$. Since $d\eta=d(\eta^{*}\tau)=\eta^{*}d\tau=-\eta^{*}\omega$, we obtain that $\eta\in\Omega^{1}(M)$ is closed if and only if $\eta:M\rightarrow T^{*}M$ is a Lagrangian immersion.
@@
    $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$;

### v3

# On Geometry and Topology of Hessian Manifolds

@@ Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is termed an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is called a Hessian metric on $(M,\nabla)$, if $g$ can be expressed locally as the Hessian quadratic form $\nabla(df)$ of some smooth potential $f$. An affine manifold equipped with a Hessian metric is termed a Hessian manifold. We study the topology and geometry of Hessian manifolds.
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and infinite torsion-freeness of its fundamental group. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the canonical and obstruction bundles, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus. Curvature properties are then discussed. As an application, we prove that all compact Hessian-Einstein manifolds are flat.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian embeddings, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature, and classify all complete Hessian surfaces.
@@
Hessian metric, Affine differential geometry, regular convex domain, Flat line bundle, Legendre transform, Lagrangian embedding, para-Kähler manifold.
@@ 1.1 Introduction and Background
A central achievement of our work is to reveal that Hessian metrics of Koszul-type are part of a much larger network of connections than previously understood. Whilst the work of H.Shima [15] established an important equivalence between the existence of a Koszul type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Corollary2.49) and admitting a space-like Lagrangian embedding (Corollary3.30) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
@@
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem2.43), revealing a remarkably rigid underlying structure, particularly in low dimensions (Corollary2.44). This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition3.56). As a important application, we then prove that all compact Hessian-Einstein manifolds are flat (Theorem3.60). An interesting geometric uncertainty principle is also established on Hessian manifolds, revealing links to quantum mechanics.
@@
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem4.5). We also derive severe constraints on the intersection form for Hessian $4$-manifolds, and then prove that any Einstein metric on a closed Hessian $4$-manifold must be flat (Proposition4.21). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
@@
This paper is organized as follows: Section1 reviews fundamental definitions and introduces a warped product construction as first examples. Section2 suggests key topological obstructions, concerning the vanishing of Euler characteristic and the infinitude of fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section3 establishes the generalised Legendre transform, and a central equivalence between Hessian metrics of Koszul type and several other geometric structures, including positive flat line bundles and space-like Lagrangian embeddings. Curvature properties and link to complex analysis are also discussed. Finally, Section4 applies these results to low dimensions, yielding a complete topological classification of Hessian surfaces and deriving strong constraints for Hessian 4-manifolds. An appendix is included to support our classification of compact Hessian manifolds with Abelian fundamental groups.
@@ 1.2 Affine and Hessian Manifolds
Throughout this article, all smooth manifolds are tacitly assumed to be connected and without boundary.
@@
Moreover, since $\nabla$ is always assumed to be torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, in particular, on an affine manifold, all parallel 1-forms are closed.
@@ 1.3 Warped Product Construction
Now we shall state the fundamental construction theorem of Hessian warped products.
@@ 2.2 Fundamental Group and Euler Number
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n$. If the fundamental group of $M$ is Abelian, then $M$ is homeomorphic to the $n$-torus $\mathbb{T}^{n}$.
@@ 2.3.1 The Picard group
Since the connection $D$ is flat, its connection 1-forms are locally exact. without loss of generality, we may assume that there exists $f\in C^{\infty}(U_{\alpha})$ such that the restriction of $D$ on $U_{\alpha}$ is $d+df$.
@@
Now we introduce the isomorphism classes in the category of flat $\mathbb{R}$-line bundle over a fixed differentiable manifold $M$.
@@ 2.3.2 The canonical line bundle
For an orientable Riemannian manifold $(M,h)$, we denote by $\operatorname{vol}_{h}$ its volume form.
@@
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
@@
Let $\eta\in\Omega^{1}(M)$ be a $1$-form potential of $((M,\nabla),h)$. View $\eta\colon M\rightarrow T^{*}M$ as an embedding. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
@@
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected fiber.
@@
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem2.38. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition2.37, we have that $\varkappa$ is nowhere vanishing by parallel transport.
Since $\mathrm{H}^{1}(M;\mathbb{Q})$ is dense in $\mathrm{H}^{1}(M;\mathbb{R})$, there exists a closed $1$-form $\hat{\varkappa}\in\Omega^{1}(M)$ such that $\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Q})$ and $\hat{\varkappa}$ is also nowhere vanishing. Define
$m:=\operatorname{min}\{k\in\mathbb{R}:k>0,k[\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Z})\}.$
Fix a base point $q$ on $M$, and define the smooth mapping $\varphi\colon M\rightarrow\mathbb{S}^{1}$ via
[…]
where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve with $\gamma(0)=q$ and $\gamma(1)=p$. Since the periods of $m[\hat{\varkappa}$] are integers, the map $\varphi$ is well-defined. Moreover, since $\hat{\varkappa}$ is nowhere vanishing, we have that $\varphi$ is a submersion. By Ehresmann’s fibration theorem, we have that $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a fiber bundle. Denote by $F$ the fiber of $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
Identifying each homotopy class in $\pi_{1}(\mathbb{S}^{1})$ with its winding number, it is readily seen that the induced homomorphism $\varphi_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(\mathbb{S}^{1})=\mathbb{Z}$ satisfies
$\varphi_{*}([\gamma])=\int_{\gamma}m\hat{\varkappa}$
for all $\gamma]\in\pi_{1}(M)$. Since $\hat{\varkappa}\neq 0$, we have that $\operatorname{Im}(\varphi_{*})\neq 0$. Therefore, there exists a positive integer $d\in\mathbb{Z}$ such that $\operatorname{Im}(\varphi_{*})=d\mathbb{Z}$. But by the minimality of the factor $m$, we must have that $d=1$ and hence $\operatorname{Im}(\varphi_{*})=\mathbb{Z}$. Now, since $M$ is connected and $\varphi_{*}$ is surjective, the long exact sequence
$\cdots\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(\mathbb{S}^{1})\longrightarrow\pi_{0}(F)\longrightarrow\pi_{0}(M)\longrightarrow\cdots$
of homotopy groups yields that $F$ is also a connected topological space. ∎
Corollary 2.44.
Let $((M,\nabla),h)$ be an closed orientable Hessian $3$-manifold. Then either $(M,h)$ is a flat Riemannian manifold, or $M$ is a $\Sigma_{g}$-bundle over $\mathbb{S}^{1}$, where $\Sigma_{g}$ is the closed orientable surface of genus $g\geq 1$.
Proof.
Assume $(M,h)$ is not flat. Then by Theorem2.43, $M$ is a $\Sigma$-bundle over $\mathbb{S}^{1}$ for some closed surface $\Sigma$. Since $\pi_{1}(M)$ is torsion-free by Proposition2.19 and $\pi_{1}(\Sigma)$ is a subgroup of $\pi_{1}(M)$ by the long exact sequence
$\cdots\longrightarrow\pi_{2}(\mathbb{S}^{1})\longrightarrow\pi_{1}(\Sigma)\longrightarrow\pi_{1}(M)\longrightarrow\cdots$
of homotopy groups, we have that $\pi_{1}(\Sigma)$ is torsion-free, and hence $\Sigma$ is orientable. Since $M$ is aspherical by Corollary2.14, we obtain that $\Sigma$ ditto. Therefore, in particular, $\Sigma$ is not homeomorphic to $\mathbb{S}^{2}$. This concludes the proof. ∎
@@ 3.1 Digression on Complex Geometry
The following totally real embedding theorem for complete Hessian manifolds is a direct application of Proposition3.3.
@@
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
@@ 3.2 Legendre Transformation
The following propositions characterize the universal coverings of compact radiant affine manifolds as open cones in the Euclidean spaces.
@@
In [25], D.Fried et al. studied the existence of parallel tensors on compact radiant affine manifolds.
Theorem 3.12 (Fried-Goldman-Hirsch,1981[25]).
A compact radiant affine manifold $(M,\nabla)$ does not admit a nonzero parallel $1$-form.
Theorem 3.13 (Fried-Goldman-Hirsch,1981[25]).
A compact radiant affine manifold $(M,\nabla)$ of dimension $n$ does not admit a nonzero parallel $n$-form.
Adapting the theorems above to our framework immediately yields the following corollaries.
Corollary 3.14.
Let $(M,\nabla)$ be a compact radiant manifold. Then the Euler vector field of $(M,\nabla)$ is unique.
Proof.
Combine Theorem3.12 with Remark3.7. ∎
Corollary 3.15.
Let $(M,\nabla)$ be a compact radiant manifold. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is non-trivial.
Proof.
Combine Theorem3.13 with Corollary2.32. ∎
Proposition 3.16.
Let $((M,\nabla),h)$ be a compact radiant Hessian manifold. Then the Riemannian manifold $(M,h)$ is not flat.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable. Now the desired result follows immediately from Theorem3.13 and Theorem2.38. ∎
@@
See [11] for a proof of the fact that the conjugate connection $D^{*}$ is well-defined and that $D^{**}=D$ for every connection $D$ on the tangent bundle $TM\rightarrow M$.
@@
3.3 Space-like Lagrangian Embeddings
In this section, we characterize Hessian manifolds of Koszul type as space-like Lagrangians, and then study their minimality as submanifolds of their own cotangent bundles.
@@ 3.3 Space-like Lagrangian Embeddings
Note that, the Riemann extension of a flat torsion-free connection is flat of neutral signature.
@@
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$, a $1$-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian embedding of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an embedded Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the Riemann extension of $\nabla$.
Remark 3.28.
Let $M$ be a differentiable manifold, and $\nabla$ a torsion-free connection on the tangent bundle $TM\rightarrow M$. Then any space-like Lagrangian embedding of $(M,\nabla)$ is by definition an exact Lagrangian, and hence admits a vanishing Maslov class $\mu\in\mathrm{H}^{2}(M;\mathbb{Z})$.
The following theorem shows that, in the category of affine manifolds, space-like embedded Lagrangians are precisely those Hessian manifolds of Koszul type.
@@
    $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$;
@@
Recall from differential topology that a smooth cross-section of a fiber bundle is always an embedding.
Denote by $\tau$ the tautological 1-form of $T^{*}M$, and denote by $\omega$ the standard symplectic form on $T^{*}M$. Then we have $\omega=-d\tau$ and $\eta^{*}\tau=\eta$. Since $d\eta=d(\eta^{*}\tau)=\eta^{*}d\tau=-\eta^{*}\omega$, we obtain that $\eta\in\Omega^{1}(M)$ is closed if and only if $\eta:M\rightarrow T^{*}M$ is a Lagrangian embedding.
@@
    $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$;